40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Jul 3 16:28:38 2014, Nichin, Update, Electronics, RF cables installed AS55_03-07-2014_162634.pdf
    Reply  Mon Jul 7 10:35:43 2014, Jenne, Update, Electronics, RF PDs needed 
       Reply  Mon Jul 7 17:20:09 2014, Nichin, Update, Electronics, RF PDs needed 
    Reply  Mon Jul 7 16:52:02 2014, manasa, Update, Electronics, RF cables need to be rerouted 
       Reply  Mon Jul 7 18:38:27 2014, Nichin, Update, Electronics, RF cables rerouted 1.png2.png
          Reply  Mon Jul 7 22:07:29 2014, rana, Update, Electronics, RF cables rerouted Untitled.pdf
          Reply  Tue Jul 15 12:36:17 2014, Nichin, Update, Electronics, RF cables rerouted 
Message ID: 10147     Entry time: Mon Jul 7 22:07:29 2014     In reply to: 10145
Author: rana 
Type: Update 
Category: Electronics 
Subject: RF cables rerouted 

This Red Ladder movement also probably included some bumping of the MC2 stack (be more careful when working in the lab). Around 5 PM today, the MC lost lock.

When I came in later, I found that the MC was flashing the TEM17,9 mode and had been misaligned like this for ~1 hour. I looked through the MC SUS sensor trends and moved MC2 back to its old position to get the locking back. I disabled the WFS, tweaked the TEM00 mode alignment using only MC2 sliders, and then re-enabled WFS. Its been OK for the past 5 hours.

Attachment 1: Untitled.pdf  251 kB  | Hide | Hide all
Untitled.pdf
ELOG V3.1.3-